OpenStudy (anonymous):

7log y-2log x+ 1/2log z express as a single logarithm and simplify if possible is the answer log y^7 (log z^1/2) ???? _________ log x^2

OpenStudy (turingtest):

Not quite. Your expression is in terms of three logs, not one. We want everything under one log sign. The three rules are\[a \log x=\log (x^a)\](You used this rule properly, by the way.) But pay close attention to the next two.\[\log a+\log b=\log(ab)\]\[\log a-\log b=\log({a \over b})\]Notice how the log signs combine. Here's what you did:\[\log a+\log b=(\log a)(\log b)\]and\[\log a-\log b={\log a \over \log b}\]These rules are wrong. Try again using the correct rules, your answer is very close!

OpenStudy (anonymous):

is it log y^7 (log z^1/2) _____ x^2

OpenStudy (turingtest):

even closer! but we still have two log signs... I will show you the answer to this one because you are so close, try to follow along and find your mistake...

OpenStudy (turingtest):

\[7\log y-2\log x+ {1\over2}\log z=\log(y^7)-\log(x^2)+\log(z^{1/2})\]\[=\log({y^7\over x^2})+\log(z^{1/2})=\log({y^7z^{1/2}\over x^2})\]you just forgot to get the z into the log sign. Like I said, look for exactly where you made your mistakes.

OpenStudy (anonymous):

wow thanks you made this much easier to see with your hints. i appreciate you taking the time to explain this to me. and I should of multiplied the z^1/2 yo just the y not the whole divisional problem:)

OpenStudy (turingtest):

thanks for listening :)